OpenStudy (anonymous):

determine whether the function y=tan x/x^2 is even, odd, or neither

OpenStudy (anonymous):

is it \[\tan \frac{ x }{ x^2}\] or \[\frac{\tan x }{ x^2}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so plug in -x for x and what do you get?

OpenStudy (anonymous):

A golfer’s score of 5 under par is represented by –5 on a number line. How would a golfer's score of 5 over par best be represented? A. –5 B. 0 C. +5 D. +10

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

Please erase and post in the regular spot

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

can anyone help

OpenStudy (anonymous):

open your own question, @africa3

OpenStudy (anonymous):

not you @bernardod , @africa3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I put -x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

even functions have\[f(-x)=f(x)\] whereas odd functions have \[f(-x)=-f(x)\]

OpenStudy (anonymous):

so when you take f(-x), does your function look the same as the original function, a negative version of the original function, or neither?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so \[f(x) = \frac{ \tan x }{ x^2 }\Rightarrow f(-x) = \frac{ \tan (-x) }{(-x)^2 }=\frac{ -\tan x }{ x^2 }=-\frac{ \tan x }{ x^2 }\]

OpenStudy (anonymous):

ooooo...I see

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

learn to plug it in and work it out to see what you get! if you plug in -x for x and get f(x) then it's and even function. if you plug in -x and get -f(x) then it's an odd funcction. if you don't get either or those then it's neither even nor odd.

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

i do watever i want ot do